MINUTES OF THE PS-ISOLDE
COORDINATION MEETING held on 24
Butler, R. Catherall, B. Crepieux, A. Dorsival, V. Fedosseev,U. Georg, U.
Koester, G. Le Scornet, M. Lindroos, S. Meunier, T. Nilsson, J. Parra-Lopez, F. Wenander
ISOLDE at URL: http://isolde.cern.ch
Minutes at URL: http://ps-div.web.cern.ch/ps-div/Meetings/ISOLDE/TechMtgs.html
Consolidation projects at URL: http://isolde-consolidation.web.cern.ch/isolde-consolidation/
PSMB meetings at URL: http://cern.web.cern.ch/CERN/Divisions/PS/Meetings/PSMB.html
The HRS was initially running and was later in the week
cooling down for a target change on Monday morning. A small collection was
done during the cooling period, which caused some questions marks concerning
the timing of the target change (UC targets -> 3 days cooling time). The
collection proved to be minor and of no importance for the cooling down but
remember to take this into account when modifying the schedule "on the
"false" vacuum interlock signal early on Monday morning caused the
control of the power supplies to block. Such a fault requires a global reset
on the local PLC.
tests on the ZrO2 target showed slightly longer release times for Ga than at
the earlier target test. still, the resulting yield was sufficient for
hot-link task in the "passerelle" for the ISOLDE control, system
went down. The result was that all "knobs" on the PC consoles were
empty. The task could be restarted by Frank Locci from his office.
problem with the control database caused the wrong values to be displayed on
the windows knobs. The database was quickly corrected by the control group.
- The control group needs urgently to have access to the tape station for
tests of the controls. Mats will organize this on Monday morning (30/9).
- A confusing message on the GPS.MAG70 control knob will be removed by the
4. Radiation protection
- Part of the target laboratory in building 179 has know been transformed to
a class B laboratory (Swiss legislation term). The area will be used to
prepare the ntof targets and especially adapted to their needs. Contact
Richard for further information if you would have a need to use such an
- Remember to always inform TIS about ANY collections made at ISOLDE. Even
if it is a small "insignificant" collection, please, discuss it
with TIS in advance. They are there to advice and to protect us.
- Remember that all staff, including industrial representants for
commissioning work, need a film badge if they are to work in the hall.
- You can submit abstracts
on the SAFERIB web site (http://cern.ch/saferib).
A second announcement has just been published on the web site including a
list of speakers. The program will be made available ASAP on the web site.
- The ST division has found a solution to make the service on the cooling
water and ventilation system at ISODLE in week 48. There will be NO cooling
water or ventilation in that week at ISOLDE. The separators will be
re-started for off-line work and separator courses in week 49.
- A separator course will organized before Christmas at ISOLDE. Please
contact Thomas Nilsson and Uwe Georg for details.
- There will be a power cut(s) 9-10/11 in the ISOLDE area. Please contact
Thomas if you need "safe power" during this period.
- A warm welcome to the new ISOLDE physics group leader, Peter Butler.
- For major installation work we need a good forewarning (several months).
All such work has to take place in the shutdown to not disturb the running
- All installations in the hall has to be inspected by TIS for safety.
Contact Uwe Georg for details.
- We note that it is difficult for experiments to get the alignment service
they need. The CERN service is doing there absolute best but they are heavily
charged and cannot make unplanned interventions with short notice. We are, together
with this group, looking for a better solution.
- The cleaning service for the hall is not working as it should. Uwe Georg
is investigating why.
1 October 2002 at 11.00
in the EP Conference Room A, building 304-1-001.
by Mats Lindroos